OpenStudy (anonymous):

Which shows this equation in standard form? 7 – 3(x – y) = 5x + 2 A. 3x – 8y = 5 B. 8x + 3y = 5 C. 8x – 3y = 5 D. 8x – 3y = –5

OpenStudy (mathmale):

The key to "solving" this problem lies in knowing what a linear equation in standard form looks like: Ax+By=C. So, if we have 7 – 3(x – y) = 5x + 2, we'll need to perform the indicated multiplication first, and then combine like terms, and, finally, to rearrange the terms with the variable x and its coefficient first, the var. y and its coeff. second, and the constant on the right of the equation. Mind starting by multiplying out that -3(x-y) in 7 – 3(x – y) = 5x + 2 ?

OpenStudy (anonymous):

ok. then what.

OpenStudy (anonymous):

why

OpenStudy (anonymous):

oh. HE wanted me to answer it. Ok. I didnt get that. Thanks

OpenStudy (anonymous):

thnaks

OpenStudy (mathmale):

Yes, I'm always looking for ways to help OpenStudy participants learn to find their own answers, with guidance and positive reinforcement. 7 – 3(x – y) = 5x + 2 should be "expanded," which, as I said before, involves multiplying out -3(x-y). -3(x-y) = - 3x + 3y So now you have 7 – 3x + 3y = 5x + 2 Please, @gymnast13, show your stuff here: Combine the like terms. When you've done that, rearrange your expression to look like Ax + By = C.